# Quantitative assessment of biological dynamics with aggregate data

### v1

@@ 4 Discussion
We present a unified statistical learning approach to parameter estimation of IVPs probed with aggregated data. Our framework reproduces the missing measurements and explicitly accounts for data aggregation, this way allowing for high modeling fidelity and improved estimation. For this task, we derived and applied specialized MCMC methods, such as cHMC and mESS, that included a modified HMC algorithm to handle summary statistics constraints in the batch data we model, as well as a modified elliptical slice sampler to navigate parameter sampling [5, 48, 47, 10].
@@
This study focuses on the analysis of biological data derived from replication measurements as commonly used in biology research. We demonstrated that our methods successfully analyze data from laboratory batch growth experiments conducted on Prochlorococcus populations and a specialized dynamical system representing the underlying behavior. Nevertheless, the generality of our method allows for applications involving diverse data sets or dynamical models.
@@
Due to its robustness, our methods can contribute to principled learning and assimilation efforts in IVPs where the underlying measurements are hidden, missing, or distorted by non-Gaussian noise and aggregation. These features are critical for extracting reliable insights from historical data sets or data sets that are confidential or obfuscated, and only down-sampled data or summary statistics are available. Using our advanced learning approach, our methods provide modelers and practitioners with enhanced tools to decipher incomplete data, ensuring the validity of their analysis despite the lack of pre-aggregated information. Our approach not only strengthens the accuracy of the inference that can be drawn, but also facilitates broader applications in the life sciences [61, 25, 39].
@@
Our novel framework leverages the Bayesian paradigm to unify all aspects of modeling biological dynamics under uncertainty within a single posterior distribution [51]. Our comprehensive approach allows simultaneous parameter estimation, recovery of lost information, and interpretation in physical terms; however, its improved performance comes with its drawbacks. In particular, our framework entails a complex statistical learning approach that requires an extensive mathematical background and advanced computational procedures that pose a barrier to practitioners. In addition, successful implementation of our framework cannot avoid costly algorithms, such as the numerical integration of the underlying ODE, or, similar to all methods based on MCMC, repetitive generation of random variates leading to generally long runs even under the efficient sampling schemes developed herewith. Characteristically, the generation of the results in each analysis of this study requires runs of $\approx 10$ hr on a typical single-core laptop computer. Reducing such a high computational time is the focus of future research that may consider specialized algorithmic improvements within problem-specific domains.
@@ Acknowledgments
This work is based on raw data provided by Erik Zinser through experimental work supported by the NSF OCE-2023680 grant.
@@ Appendix A Summary of the framework’s equations
where $\mathbb{B},\mathbb{C}$ are prior distributions of the dynamical $g$ and observation $h$ parameters, $\mathbb{A}$ is the likelihood of the raw measurements $y_{1:N}^{1:M}$, $F(\cdot)$ is the observation function, and $x^{g}(\cdot)$ is the solution of the underlying IVP. In this description, the aggregated data is $z_{1:N}^{1:M}$, the evaluation times are $t_{1:N}$, the statistics are $G^{1:M}(\cdot)$, and the batch size is $K$.
@@ Appendix B Definitions of probability distributions
$\delta_{a}(E)=\begin{cases}1&\text{if }a=E\\ 0&\text{if }a\neq E\end{cases}$
@@ C.1 Multiplicative elliptical slice sampling
Input: Current state $\mathbf{g}$, auxiliary variance $\sigma^{2}$, likelihood function $L$
@@
$\mathbf{\nu}\sim\mathcal{N}(0,S)$triangleright Choose ellipse
$\mathbf{\lambda}\sim\mathcal{N}(0,\sigma^{2})$triangleright Choose auxiliary variable
@@ C.2 Standard Hamiltonian Monte Carlo
where $M$ is a “mass” matrix. This choice for the Hamiltonian function leverages the known gradient information of our target distribution $\pi(q)$ [5, 8]. We can convert our proportionality relationship
@@
$p^{(j)}\sim\mathcal{N}(0,M),$ (31)
where $\mathcal{N}(0,M)$ is a multivariate normal. For the Metropolis step, we compute the acceptance ratio $R^{(j)}$ using the expression
@@
As in the Metropolis algorithm, we accept the proposed sample and set $q^{(j+1)}=\tilde{q}^{(j+1)}$ with probability $\min(1,R^{(j)})$, otherwise we set $q^{(j+1)}=q^{(j)}$. We summarize the algorithm below.
@@
If the integration is exact or if a reversible, measure-preserving numerical integrator (such as Störmer-Verlet) is used, then this scheme satisfies a detailed balance [5], so the empirical statistics of the resulting Markov chain $\{q^{(j)}\}_{j=0:J}$ approach those of the target distribution with density $\pi(q)$.
In practice the exact integration of the Hamiltonian is not possible, but the better we can approximately integrate the Hamiltonian, the higher our acceptance rate will be [8, 5]. The standard numerical integrator for HMC is the explicit, second order Störmer-Verlet numerical scheme [37, 8, 5]. As it is the base algorithm from which RATTLE is derived, we give one step (step size $h$) of the Störmer-Verlet numerical scheme here.
@@
The Störmer-Verlet scheme is indeed reversible and preserves the phase-space measure, so the resulting Markov chain has the desired approximation properties.
@@ C.3 Constrained Hamiltonian Monte Carlo
In this section, we follow the notation used in the previous section and outline the constrained Hamiltonian Monte Carlo algorithm as described by Brubaker et al. [9]. The goal of cHMC is to sample a surface measure, in particular a probability distribution with density $\pi(q)\propto f(q)$ known only up to a multiplicative constant supported on a differentiable manifold $\mathcal{M}$. We take $\mathcal{M}$ to be embedded in Euclidean space and defined as the zero set of a constraint function $q\mapsto c(q)\in\mathbb{R}^{d}$ with full-rank Jacobian $J_{c(q)}$ for all $q\in\mathcal{M}$. We will use the notation $\nabla c(q)=J_{c(q)}^{T}$ in analogy to the gradient for the transpose of the Jacobian. As in the unconstrained setting, we introduce a virtual momentum variable $p$.
To represent data aggregation, the set of values $q=y_{1:N}^{1:K}$ which simultaneously satisfy eq. 1 is parametrized as the solution set of the following $2N$ constraint functions, $G^{1}(y_{n}^{1:K})=z_{n}^{1}$ and $G^{2}(y_{n}^{1:K})=z_{n}^{2}$ for $n=1:N$, where $z_{n}^{1}$ and $z_{n}^{2}$ are the mean and variance, respectively, of the $K$ measurements taken at time $t_{n}$. Here, $G^{1}$ and $G^{2}$ are as defined in eqs. 2 and 3. It is convenient for the implementation to reparametrize these $2N$ constraints as follows. We constrain samples $q=y_{1:N}^{1:K}$ to lie on the zero locus of
$c_{n}^{1}(q)$ $=\frac{1}{K}\sum_{k=1}^{K}y_{n}^{k}-z_{n}^{1},$ $n$ $=1:N,$ (33)
$c_{n}^{2}(q)$ $=\frac{1}{K}\sum_{k=1}^{K}(y_{n}^{k})^{2}-\tilde{z}_{n}^{2}$ $n$ $=1:N,$ (34)
where $\tilde{z}_{n}^{2}$ is the raw (uncentered) second moment of the $K$ measurements recorded at time $t_{n}$. We collect all constraints into a column vector-valued function $c=(c_{1}^{1},\cdots,c_{N}^{1},c_{1}^{2},\cdots,c_{N}^{2})^{T}$ which reduces the satisfaction of the $2N$ mean and variance constraints to the simple expression $c(q)=0$.
@@
Let $T_{q}\mathcal{M}$ denote the plane tangent to $\mathcal{M}$ at $q$, and let $\mathcal{N}^{(j)}$ denote the surface measure in $T_{q^{(j)}}\mathcal{M}$ induced by the normal distribution $\mathcal{N}(0,M)$ in the ambient Euclidean space. We take the initial momentum sample $p^{(j)}$ from the singular distribution $\mathcal{N}^{(j)}$. Under these initial conditions, the solution $(q,p)$ remains in the tangent bundle $T\mathcal{M}=\sqcup_{q\in\mathcal{M}}T_{q}\mathcal{M}$ of $\mathcal{M}$ [19, 33, 34].
@@
The integration can be exact or approximate. We compute the Metropolis acceptance ratio $R^{(j)}$ using the same expression as in the unconstrained case. With these particulars, we can give the cHMC algorithm.
@@
Since a numerical integrator with energy-conserving properties that respects the constraint structure is desired, the standard choice for a numerical integrator is the second-order method RATTLE [9, 35, 28]. In the context of molecular dynamics, it has been shown that RATTLE is reversible and preserves Lebesgue surface measure in $T\mathcal{M}$ (see [33, 34]), and hence satisfies a detailed balance. So, the Markov chain generated by cHMC has the desired statistical approximation properties. Here we describe one step of RATTLE with step size $h$.
@@
Though based on Störmer-Verlet, the RATTLE scheme is quite computationally distinct. Each step of RATTLE requires, in general, the solving of a nonlinear and a linear system to find the appropriate multipliers $\lambda_{(q)}$ and $\lambda_{(p)}$, respectively. We implement these solvers with the Newton-Raphson method. The condition on $\lambda_{(q)}$ keeps $\tilde{q}\in\mathcal{M}$ while $\lambda_{(p)}$ keeps $\tilde{p}\in T_{q}\mathcal{M}$ ensuring that the output $(\tilde{q},\tilde{p})$ remains in $T\mathcal{M}$.
@@ Appendix D Least squares parameter estimation
Here, $x^{g}(t_{n})$ denotes the solution of the IVP in eq. 5 with parameters $g$ evaluated at time $t_{n}$, $F(\cdot)$ is the observation function in eq. 4, and $z_{n}$ denotes the summary statistic computed in eq. 1.

### v2

@@ 4 Discussion
In this study, we present a unified statistical learning approach to the estimation of the parameters of IVPs probed with aggregated data. Our framework reproduces the missing measurements and explicitly accounts for data aggregation, this way allowing for high modeling fidelity and improved estimation. For our learning tasks, we derived and applied specialized MCMC methods, such as cHMC and mESS, that include a modified HMC sampler to handle summary statistics constraints in the data we model, as well as a modified elliptical slice sampler to navigate parameter sampling [5, 58, 57, 11]. Developed within a fully Bayesian methodology, our framework can be readily used to provide point estimates of the parameters of interest as well as uncertainty quantification.
@@
Our study focuses on the analysis of biological data derived from replication measurements as commonly used in biological research. We demonstrated that our methods successfully analyze data from laboratory batch growth experiments conducted on Prochlorococcus cultures and a specialized dynamical system representing the underlying behavior. Nevertheless, the generality of our method allows for applications involving diverse data sets or dynamical models beyond those tested in this study.
@@
Due to its robustness, our methods can contribute to principled learning and assimilation efforts in IVPs where the underlying measurements are hidden, missing, or distorted by non-Gaussian non-additive noise and aggregation. These features are critical for extracting reliable insights from historical data sets or data sets that are confidential or obfuscated, and only down-sampled data or summary statistics are available. Using our advanced learning approach, our methods provide modelers and practitioners with enhanced tools to decipher incomplete data, ensuring the validity of their analysis despite the lack of pre-aggregated information. Our approach not only strengthens the accuracy of the inference that can be drawn, but also facilitates broader applications in the biological and life sciences [82, 29, 47].
@@
Our novel framework leverages the Bayesian paradigm to unify all aspects of modeling biological dynamics under uncertainty within a single posterior distribution [61]. Our comprehensive approach allows simultaneous parameter estimation, recovery of lost information, and interpretation in physical terms; however, its improved performance comes with its drawbacks. In particular, our framework entails a complex statistical learning approach that requires an extensive mathematical background and advanced computational procedures that pose a barrier to practitioners. To help a wider adoption of the developed methods, we provide a prototype computational implementation in [51]. Our implementation is developed in the Matlab programming language [28] and solves the synthetic Prochlorococcus scenarios of sections 2.2 and 3.1.
Nevertheless, a successful implementation of our framework cannot avoid costly algorithmic steps, such as the numerical integration of the underlying ODE, or, similar to all methods based on MCMC, repetitive generation of random variates leading to generally long runs even under the efficient sampling schemes developed herewith. Characteristically, the generation of the results in each analysis of this study with our prototype implementation requires runs of $\approx 5$ min, on a typical single-core laptop computer (Apple MacBook Pro, 2022 model). This time is contrasted with least squares fitting that is completed within $\approx 1$ min. Reducing such a high computational time is the focus of future research that may consider specialized algorithmic solutions within problem-specific domains.
@@ Acknowledgments
This research is based on raw data provided by Erik Zinser through laboratory work supported by the NSF OCE-2023680 grant.
@@ Appendix A Summary of the framework’s equations
where $\mathbb{B},\mathbb{C}$ are the prior distributions of the dynamical $g$ and observation $h$ parameters, $\mathbb{A}$ is the likelihood of the raw measurements $y_{1:N}^{1:M}$, $F(\cdot)$ is the observation function, and $x^{g}(\cdot)$ is the solution of the underlying IVP with initial conditions applied at time $t_{0}$. In this description, the aggregated data is $z_{1:N}^{1:M}$, the evaluation times are $t_{1:N}$, the statistics are $G^{1:M}(\cdot)$, and the batch size is $K$.
@@ Appendix B Definitions of probability distributions
$\delta_{a}(E)=\begin{cases}1&\text{if }a=E\\ 0&\text{if }a\neq E\end{cases}.$
•  The Bernoulli probability mass function is
$\text{Bernoulli}(k;p)=\begin{cases}p&\text{if }k=1,\\ 1-p&\text{if }k=0\\ 0&\text{otherwise}\end{cases}.$
•  The Uniform probability density on the interval $(a,b)$ is
$\text{Uniform}_{(a,b)}(x)=\begin{cases}\frac{1}{b-a}&\text{if }x\in(a,b)\\ 0&\text{otherwise}\end{cases}.$
•  The $d$-dimensional Normal probability density is
$\text{Normal}(x;\mu,\Sigma)=\frac{1}{\sqrt{(2\pi)^{d}|\Sigma|}}\exp\left(-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)\right).$
@@ C.1 Multiplicative elliptical slice sampling
Input: Current state $\mathbf{g}$, auxiliary variance $\sigma^{2}$, likelihood function $L(\cdot)$
@@
$\mathbf{\nu}\sim\text{Normal}(0,S)$triangleright Choose ellipse
$\mathbf{\lambda}\sim\text{Normal}(0,\sigma^{2})$triangleright Choose auxiliary variable
@@ C.2 Standard Hamiltonian Monte Carlo
where $M$ is a “mass” matrix. This choice for the Hamiltonian function [5, 9] leverages the known gradient information of our target distribution $\pi(q)$. We can convert our proportionality relationship
@@
$p^{(j)}\sim\text{Normal}(0,M),$ (31)
where $\text{Normal}(0,M)$ is a multivariate normal. For the Metropolis step, we compute the acceptance ratio $R^{(j)}$ using the expression
@@
As in the Metropolis algorithm, we accept the proposed sample and set $q^{(j+1)}=\tilde{q}^{(j+1)}$ with probability $\min(1,R^{(j)})$, otherwise we set $q^{(j+1)}=q^{(j)}$. We summarize the entire scheme in algorithm 2.
@@
Provided the integration is exact or a reversible, measure-preserving numerical integrator (such as Störmer-Verlet) is used, then this scheme satisfies detailed balance [5], so the statistics of the resulting Markov chain $\{q^{(j)}\}_{j=0:J}$ approach those of the target distribution with density $\pi(q)$.
In practice the exact integration of the Hamiltonian is not possible; nevertheless, better approximate integration of the Hamiltonian leads to higher acceptance rates [9, 5]. The standard numerical integrator for HMC is the explicit, second order Störmer-Verlet numerical scheme [45, 9, 5]. As it is the base algorithm from which RATTLE is derived, we give one step (step size $h$) of the Störmer-Verlet numerical scheme in algorithm 3.
@@
The Störmer-Verlet scheme is indeed reversible and preserves the phase-space measure [9, 5], so the resulting Markov chain has the desired approximation properties.
@@ C.3 Constrained Hamiltonian Monte Carlo
In this section, we follow the notation used in section C.2 and outline the constrained Hamiltonian Monte Carlo algorithm as described by Brubaker et al. [10]. The goal of cHMC is to sample a surface measure, in particular a probability distribution with density $\pi(q)\propto f(q)$ known only up to a multiplicative constant supported on a differentiable manifold $\mathcal{M}$. We take $\mathcal{M}$ to be embedded in Euclidean space and defined as the zero set of a constraint function $q\mapsto c(q)\in\mathbb{R}^{d}$ with full-rank Jacobian $J_{c(q)}$ for all $q\in\mathcal{M}$. We will use the notation $\nabla c(q)=J_{c(q)}^{T}$ in analogy to the gradient for the transpose of the Jacobian. As in the unconstrained setting, we introduce a virtual momentum variable $p$.
To represent data aggregation, the set of values $q=y_{1:N}^{1:K}$ which simultaneously satisfy eq. 1 is parametrized as the solution set of the following $2N$ constraint functions, $G^{1}(y_{n}^{1:K})=z_{n}^{1}$ and $G^{2}(y_{n}^{1:K})=z_{n}^{2}$ for $n=1,\dots,N$, where $z_{n}^{1}$ and $z_{n}^{2}$ are the mean and standard deviation, respectively, of the $K$ measurements taken at time $t_{n}$. Here, $G^{1}(\cdot)$ and $G^{2}(\cdot)$ are as defined in eqs. 2 and 3.
It is convenient for the implementation to reparametrize these $2N$ constraints as follows. We constrain samples $q=y_{1:N}^{1:K}$ to lie on the zero locus of
$c_{n}^{1}(q)$ $=\sum_{k=1}^{K}y_{n}^{k}-K\tilde{z}_{n}^{1},$ $n$ $=1,\dots,N,$ (33)
$c_{n}^{2}(q)$ $=\frac{\tilde{z}_{n}^{1}}{2}\left(\frac{1}{\tilde{z}_{n}^{2}}\sum_{k=1}^{K}(y_{n}^{k})^{2}-K\right)$ $n$ $=1,\dots,N,$ (34)
where the new constants $\tilde{z}_{n}^{1}$ and $\tilde{z}_{n}^{2}$ are given in terms of $z_{n}^{1}$ and $z_{n}^{2}$ according to
$\tilde{z}_{n}^{1}$ $=z_{n}^{1}$ $n$ $=1,\dots,N,$
$\tilde{z}_{n}^{2}$ $=(z_{n}^{1})^{2}+\frac{K-1}{K}(z_{n}^{2})^{2}$ $n$ $=1,\dots,N.$
The new constants are the batch raw (uncentered) first and second moments [61] of the measurements recorded at time $t_{n}$.
For convenience of the notation, we collect all constraints into a column vector-valued function $c=(c_{1}^{1},\cdots,c_{N}^{1},\allowbreak c_{1}^{2},\cdots,c_{N}^{2})^{T}$ which reduces the satisfaction of the $2N$ mean and standard deviation constraints to the simple expression $c(q)=0$.
@@
Let $T_{q}\mathcal{M}$ denote the plane tangent to $\mathcal{M}$ at $q$, and let $\mathcal{N}^{(j)}$ denote the surface measure in $T_{q^{(j)}}\mathcal{M}$ induced by the normal distribution $\text{Normal}(0,M)$ in the ambient Euclidean space. We take the initial momentum sample $p^{(j)}$ from the singular distribution $\mathcal{N}^{(j)}$. Under these initial conditions [22, 41, 42], the solution $(q,p)$ remains in the tangent bundle $T\mathcal{M}=\sqcup_{q\in\mathcal{M}}T_{q}\mathcal{M}$ of $\mathcal{M}$.
@@
The integration can be exact or approximate. We compute the Metropolis acceptance ratio $R^{(j)}$ using the same expression as in the unconstrained case. Our entire scheme is summarized in algorithm 4.
@@
Since a numerical integrator with energy-conserving properties that respects the constraint structure is desired, the standard choice for a numerical integrator is the second-order method RATTLE [10, 43, 36]. It has been shown that RATTLE is reversible and preserves Lebesgue surface measure in $T\mathcal{M}$, for instance see [41, 42], and hence satisfies detailed balance. So, the Markov chain generated by cHMC has the desired statistical approximation properties. In algorithm 5, we describe one step of RATTLE with step size $h$.
@@
Though based on Störmer-Verlet, the RATTLE scheme is quite computationally distinct. Each step of RATTLE requires, in general, the solving of a nonlinear and a linear system to find the appropriate multipliers $\lambda_{(q)}$ and $\lambda_{(p)}$, respectively. We implement these solvers with the Newton-Raphson method [69]. The condition on $\lambda_{(q)}$ keeps $\tilde{q}\in\mathcal{M}$ while $\lambda_{(p)}$ keeps $\tilde{p}\in T_{q}\mathcal{M}$ ensuring that the output $(\tilde{q},\tilde{p})$ remains in $T\mathcal{M}$ as desired.
@@ Appendix D Least squares parameter estimation
Here, $x^{g}(t_{n})$ denotes the solution of the IVP in eq. 5 with parameters $g$ evaluated at time $t_{n}$, $F(\cdot)$ is the observation function in eq. 4, and $z_{n}$ denotes the mean summary statistic computed in eq. 1.